OpenStudy (anonymous):

will award medal Let R be the rectangle with vertices (0,0), (2,0), (2,2), and (0,2) and let f(x,y) = \sqrt{xy}. (a) Find reasonable upper and lower bounds for \int_{R}f\,dA without subdividing R.

OpenStudy (anonymous):

@Zarkon youre smart as hell, anyway you could help me out? to at least start this problem

OpenStudy (zarkon):

depends on what they mean by reasonable. I would take \(\sqrt{0\cdot0}\) and \(\sqrt{2\cdot2}\) as lower and upper bounds for \(f\) then multiply by the area of \(R\)

OpenStudy (zarkon):

\[0\le \sqrt{xy}\le2\] so \[\iint\limits_{R}0dA\le\iint\limits_{R}\sqrt{xy}dA\le\iint\limits_{R}2dA\]

OpenStudy (zarkon):

\[0\times Area(R)\le\iint\limits_{R}\sqrt{xy}dA\le2\times Area(R)\] \[0\le\iint\limits_{R}\sqrt{xy}dA\le2\times4=8\]

OpenStudy (anonymous):

so 8 would be the upper bound

OpenStudy (zarkon):

yes...though it is not a great upper bound

OpenStudy (anonymous):

and zero is the lower bound

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

the true value of the integral is 32/9 so I guess 0 and 8 aren't terrible

OpenStudy (anonymous):

the webwork likes em. and for the second part

OpenStudy (anonymous):

for making a table what would i do to set that up?

OpenStudy (zarkon):



OpenStudy (anonymous):

so would my delta x and delta y be 2?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

take each square at a time. find the max and min of \(\sqrt{xy}\) on each square

OpenStudy (anonymous):

 would it be this? @Zarkon